Shift report 20010816 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Thu Aug 16 2001 - 08:04:23 EDT

  • Next message: Flemming Videbaek: "Thursdays meeting"

    ------------------------------------------------------------
    Date: Thu Aug 16 08:04:23 2001, Shift: 00:00-08:00
    Supervisor: Alexandru JIPA, with: Andrei MAKEEV
    Log book#7 pages 97 - 100
    
    Shift Summary:
    ------------------------------------------------------------
    1. At 0.05 - Injecting beam. From 0.40 - top energy, beam 
    cogged and stored. Since 0.55 - runs.
    2. Run 4224 with trigger 4 not downscaled (0.55-1.25).
    Run 4225 with trigger 1 only (test requested by Dr.Ramiro
    Debbe) - between 1.26 and 1.56
    Run 4226 - with triggers 1,4 and 8 (2.16-7.18). This run 
    can be affected by the presence of the ZDCs pulser (see page
    100 from LOG-BOOK #7). Pulser stopped by Andrei Makeev.
    Run 4227 - with triggers 1,4 and 8 - since 7.39; run stopped
    at 8.05 because are problems with the beam
    3. In the absence of the beam the high voltages on TPM1, TPM2,
    T1 and T2 must be modified (Values are posted on the wall -
    in colour - by Dr. Ramiro Debbe). These values was checked
    periodically during the shift. No modifications observed.  
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Thu Aug 16 2001 - 08:05:11 EDT